# Asymmetry Analysis of Bilateral Shapes

### v1

@@ 4 Analyses of the Smile Data
Our smile data contains $N=25$ subjects, of which $N_{1}=12$ are control subjects and $N_{2}=13$ are cleft subjects. There are $K=24$ landmarks on the lip periphery. Figure 3 shows the lip configuration in the $x$-$y$ plane. There are $K_{P}=11$ landmark pairs and $K_{S}=2$ solo landmarks. The landmark indices are summarized in Table 3. The $M=3$ dimensional coordinates of these landmarks have been extracted at three frames: first (closed lip), middle (middle of the smile) and last frame (maximum open lip smile). Let $g=1,2$ denote the control and cleft groups, respectively.
@@
Figure 3: Landmark indices for $x$-$y$ coordinates on the lip periphery of a control subject at first frame.
@@
Table 3: Indices of landmark pairs and solos in Figure 3 on the lip periphery.
@@
In the rest of this section several test statistics will be presented in two parts; combine-then-compare approach and compare-then-combine approach.
Combine-then-compare approach using (3.9). Patel et al. (2023) carried out standard two-sided two-sample t-tests using the statistic (3.9), with conventional t-tables used to judge significance. Here we report the results of the one-sided versions of these t-tests since based on hypotheses in (3.1), one-sided tests are more appropriate. The results of our t-tests are given in Table 4. We reach the same conclusions as in Patel et al. (2023) for their two-sided tests, namely, $H_{0}$ is rejected at all three frames. Further, the one-sided t-test indicates that there is more asymmetry for cleft lip subjects versus controls. Note that here the $p$-value is the smallest at the first frame, which suggests that the two groups are the most different at the start.
Table 4: Mean, variance, t-values and $p$-values for $\phi^{*}_{L_{1}}(\bm{a})$ (equation (3.9)) for cleft and control subjects at first, middle and last frames (Patel et al. (2023)).
Cleft Control One-sided test
mean sd mean sd $t$-values $p$-values
First 1.78 0.46 1.12 0.41 3.86 0.0004(***)
Middle 1.96 0.70 1.46 0.47 2.14 0.02(*)
Last 2.14 0.80 1.55 0.43 2.34 0.01(**)
@@
Combine-then-compare approach using (3.6) and (3.7). We use the scores $\phi_{L_{1}}(\bm{a})$ and $\phi_{L_{2}}(\bm{a})$ defined in equations (3.6) and (3.7) respectively. Instead of considering asymmetries with respect to each landmark as in (3.9), we view all features in $\bm{a}$ as a whole and take the $L_{1}$ and $L_{2}$ norms. For the subjects which are more symmetric, their corresponding $\bm{a}_{n}$ should be close to the origin. Hence, the $L_{1}$ and $L_{2}$ distances between $\bm{a}_{n}$ with $\bm{0}\in\mathbb{R}^{J}$ should be smaller for control subjects. Several one-sided two-sample Mann-Whitney U tests have been performed and the results are shown in Table 5. For the last two rows in the table, the weighted $\phi_{L_{1}}(\bm{a})$ and weighted $\phi_{L_{2}}(\bm{a})$ are used and $w_{j}$ are chosen in an adaptive way as the following:
  1. 1.
    Compute the sample mean shape on the augmented data, which contains basis registered original configurations and their reflections through midplane. Since the data has been pre-registered, the sample mean shape is simply the arithmetic mean for each landmark.
  2. 2.
    The Euclidean distance within each landmark pair for this mean shape is computed and its reciprocal is used as the weight. The weights are the same among these subjects from both groups. The weights for solo landmarks are the unit length based on the scale of data.
@@
Table 5: $p$-values from one-sided two-sample Mann-Whitney U tests using $\phi_{L_{1}}(\bm{a})$ (3.6) and $\phi_{L_{2}}(\bm{a})$ (3.7).
@@
It can be seen from Table that 5 $H_{0}$ is rejected in all cases. In summary, there are significant differences between cleft and control subjects at all frames, with control subjects less asymmetric than cleft subjects overall.
Since the test results indicate significant differences between the two groups, we would like to visualize the separation between the distributions of composite asymmetry scores of the two groups. Hence, dot plots similar to the stem leaf plot are created and given in Figures 4, 5, 6, 7 and 8. In these dot plots, composite asymmetry scores for both groups are plotted together at each frame.
Figure 4 shows the dot plots for $\phi^{*}_{L_{1}}(\bm{a})$ defined in equation (3.9), whereas Figure 5 and 6 display the dot plots of $\phi_{L_{1}}(\bm{a})$ and $\phi_{L_{2}}(\bm{a})$ (equations (3.6) and (3.7)) respectively. According to these dot plots, substantial overlaps between cleft and control subjects can be seen. Moreover, the three composite asymmetry scores used in this section all have similar performances, as the separation between the two groups at each frame revealed by the dot plots are all similar.
i First frame
ii Middle frame
iii Last frame
Figure 4: Dot plots of $\phi^{*}_{L_{1}}(\bm{a})$ defined in equation (3.9) at the three frames. The top row of the graph is for cleft lip subjects whereas the bottom row is for controls.
i First frame
ii Middle frame
iii Last frame
Figure 5: Dot plots of $\phi_{L_{1}}(\bm{a})$ (3.6) at the three frames. The top row of the graph is for cleft lip subjects whereas the bottom row is for control subjects.
i First frame
ii Middle frame
iii Last frame
Figure 6: Dot plots of $\phi_{L_{2}}(\bm{a})$ (3.7) at the three frames. The top row of the graph is for cleft lip subjects whereas the bottom row is for control subjects.
Figure 7 and 8 show the dot plots for weighted $\phi_{L_{1}}(\bm{a})$ and weighted $\phi_{L_{2}}(\bm{a})$ respectively. According to these figures, these two composite scores push outliers farther away, i.e., they are sensitive to outliers, especially for weighted $\phi_{L_{2}}(\bm{a})$. So the weighted $\phi_{L_{1}}(\bm{a})$ and weighted $\phi_{L_{2}}(\bm{a})$ emphasize which are more extreme in cleft lip group.
i First frame
ii Middle frame
iii Last frame
Figure 7: Dot plots of weighted $\phi_{L_{1}}(\bm{a})$ at the three frames. The top row of the graph is for cleft lip subjects whereas the bottom row is for control subjects.
i First frame
ii Middle frame
iii Last frame
Figure 8: Dot plots of weighted $\phi_{L_{2}}(\bm{a})$ at the three frames. The top row of the graph is for cleft lip subjects whereas the bottom row is for control subjects.
@@
Compare-then-combine approach using (3.4). The method described in Section 3.2 is used to select the landmarks which possess significant asymmetry information. $v_{j}$ is selected to be the two-sample t-statistics (equation (3.15)). $V$ defined in (3.16) is computed and its critical value at $5\%$ significance level is determined via bootstrap. The total number of bootstrap iterations used was 10000. Boyett and Shuster (1977) suggest 17000 iterations to make sure the significance level is $1\%$. However, the results of using 10000 and 17000 iterations turned out to be the same for our smile data. The central pair, landmarks 6 and 8, correspond to $t$-values which exceed these critical points at all three frames. Thus, these two landmarks have the most significant information of asymmetry. Further, the simultaneous confidence interval related to landmarks 6 and 8 excludes 0, so that $H_{0}$ is rejected when use these landmarks.
@@
Further, most of the extreme $p$-values are obtained from Mann-Whitney U tests related to landmarks 6, 8 and two solo landmarks 7 and 19. These tests are performed using $\phi_{L_{1}}$ and $\phi_{L2}$ computed on various subsets of landmarks. This suggests that the two groups are differed at most when using these four landmarks. So these landmarks can carry important information on asymmetries. This finding matches the expert’s opinion. Figure 9 shows a photo of a cleft patient, where the central pair (6, 8) and solo landmarks, 7 and 19, are marked on the photo and the landmarks are same as in Figure 3. Roughly, it is a ‘Y’-shape if we join the landmarks 6, 7, 8 and then 7 with 19 by line segments. From medical point of view, this particular shape is known a priori to be the most important for the asymmetry assessment.
@@
We also applied the Welch version of the $t$-test (Welch, 1947), which accommodates different variances for the two groups. In this example, the results of the Welch $t$-test is found to be similar to the standard $t$-test (Table 4). A possible reason is that our sample size is small, so it results into similar outcome. Further, the multivariate Hotelling’s $T^{2}$ test on unsigned elementary feature vector $\bm{a}_{n}$, $n=1,\ldots,N$, is an alternative approach to the UIT. In this example, the test results fail to improve the outcomes of the univariate tests. It could be since a Hotelling’s $T^{2}$ test treats the features separately so it does not gain any power if the features point in the same direction; one-sided Hotelling’s $T^{2}$ could be appropriate approach, which we will pursue in future. In this example, the sample size is less than the dimension of $\bm{a}_{n}$, i.e. $N<J$, so the covariance matrix is singular, which requires some modifications and will be dealt in future.
@@
We end this section by summarizing the main conclusions for this smile data:
@@
  • There are statistically significant differences between the two groups at all frames. So overall, control subjects are less asymmetric than the cleft lip subjects.
@@
  • Substantial overlaps can be found on dot plots in Figure 4, 5 and 6. This matches the expert’s expectation that some cleft lip subjects should be fairly close to normal subjects after surgeries.

### v2

@@ 5.1 Details
We now illustrate the methods using the cleft lip smile data. The data is collected using the Di4D system (Patel et al., 2023) as described in Section 2 and it has been pre-registered using the basis registration. This data contains $N=25$ subjects, of which $N_{1}=12$ are control subjects and $N_{2}=13$ are cleft subjects. There are $K=24$ landmarks on the lip periphery. Figure 3 shows the lip configuration in the $x$-$y$ plane. There are $K_{P}=11$ landmark pairs and $K_{S}=2$ solo landmarks. The landmark indices are summarized in Table 3. The $M=3$ dimensional coordinates of these landmarks have been extracted at three frames: first (closed lip), middle (middle of the smile) and last frame (maximum open lip smile). We have $J=MK_{P}+K_{S}=35$. Let $g=1,2$ denote the control and cleft groups, respectively.
@@
Figure 3: Landmark indices for $x$-$y$ coordinates on the lip periphery of a control subject at first frame from Smile Data 1.
@@
Table 3: Indices of landmark pairs and solos in Figure 3 on the lip periphery for the Smile Data 1.
@@
In the rest of this section several test statistics as described in Section 3 will be presented under two headings “combine-then-compare approach” and “compare-then-combine approach”.
5.2 Test Results for Smile Data 1
5.2.1 Application of the Combine-then-compare Approach
Patel et al. (2023) carried out standard two-sided two-sample t-tests using the statistic (3.9), with conventional t-tables used to judge significance. Here we report the results of the one-sided versions of these t-tests since based on hypotheses in (3.1), one-sided tests are more appropriate. The results of our t-tests are given in Table 4. We reach the same conclusions as in Patel et al. (2023) for their two-sided tests, namely, $H_{0}$ is rejected at all three frames. Further, the one-sided t-test indicates that there is more asymmetry for cleft lip subjects versus controls. Note that here the $p$-value is the smallest at the first frame, which suggests that the two groups are the most different at the beginning of the smile.
Table 4: Mean, variance, t-values and $p$-values for $\phi^{*}_{L_{1}}(\bm{a})$ using Smile Data 1 (equation (3.9)) for cleft and control subjects at first, middle and last frames (Patel et al. (2023)).
Cleft Control One-sided test Two-sided test
mean sd mean sd $t$-values $p$-values $p$-values
First 21.35 5.54 13.39 5.11 -3.72 0.0006(***) 0.0011(**)
Middle 23.57 8.37 17.35 5.87 -2.13 0.02(*) 0.04(*)
Last 25.65 9.66 18.22 5.21 -2.36 0.01(*) 0.03(*)
@@ 5.2.1 Application of the Combine-then-compare Approach
Then, we use the scores $\phi_{L_{1}}(\bm{a})$ and $\phi_{L_{2}}(\bm{a})$ defined in equations (3.6) and (3.7) respectively. Several one-sided two-sample Mann-Whitney U tests have been performed and the results are shown in Table 5. For the last two rows in the table, the weighted $\phi_{L_{1}}(\bm{a})$ and weighted $\phi_{L_{2}}(\bm{a})$ are used and $w_{j}$ are chosen in an adaptive way as the following:
  1. (a)
    Compute the sample mean shape on the doubled dataset, which contains basis registered original configurations and their reflections through midplane. Since the data has been pre-registered, the sample mean shape is simply the arithmetic mean for each landmark.
  2. (b)
    The Euclidean distance within each landmark pair for this mean shape is computed and its reciprocal is used as the weight. The weights are the same for all 25 subjects. The weights for solo landmarks are the unit length based on the scale of data ($w_{j}=1$ in (3.6) and (3.7)).
@@
Table 5: $p$-values from one-sided two-sample Mann-Whitney U tests using $\phi_{L_{1}}(\bm{a})$ (3.6) and $\phi_{L_{2}}(\bm{a})$ (3.7) based on Smile Data 1.
@@
It can be seen from Table 5 that $H_{0}$ is rejected in all cases. In summary, there are significant differences between cleft and control subjects at all frames, with control subjects less asymmetric than cleft subjects overall.
5.2.2 Application of the Compare-then-combine Approach
@@
Feature selection The method described in Section 3.2 is used to select the landmarks which possess significant asymmetry information. $v_{j}$ is selected to be the two-sample t-statistics (equation (3.15)). $V$ defined in (3.16) is computed and its critical value at $5\%$ significance level is determined via bootstrap. The total number of bootstrap iterations used was 10000. Boyett and Shuster (1977) has suggested 17000 iterations to make sure the significance level is $1\%$. However, the results of using 10000 and 17000 iterations turned out to be the same for our Smile Data 1. The central pair, landmarks 6 and 8, corresponds to $t$-values which exceed these critical points at all three frames. Thus, these two landmarks carry the most significant information of asymmetry.
@@
Further, most of the extreme $p$-values obtained from Mann-Whitney U tests are related to landmarks 6, 8 and two solo landmarks 7 and 19. These tests are performed using $\phi_{L_{1}}$ and $\phi_{L2}$ computed on various subsets of landmarks. This suggests that the two groups are differed at most when using these four landmarks. So these landmarks can carry important information on asymmetries. This finding matches the expert’s opinion. Figure 4 shows a photo of a cleft patient, where the central pair (6, 8) and solo landmarks, 7 and 19, are marked on the photo and the landmarks are same as in Figure 3. Roughly, it is a ‘Y’-shape if we join the landmarks 6, 7, 8 and then 7 with 19 by line segments. From medical point of view, this particular shape is known a priori to be the most important for the asymmetry assessment.
@@
Meta-analysis Fisher’s method and Pearson’s method defined in equation (4.1) and (4.2) respectively are performed. Mann-Whitney U test is used to obtain individual left-sided $p$-values, $\tilde{p}_{j}$, for $j=1,\dots,35$, using $\bm{a}_{n}$, $n=1,\dots,25$.
For Fisher’s method, we compare $P_{L}$ with the tail of $\chi_{70}^{2}$ distribution (as $J=35$) to obtain overall $p$-values. For the Pearson’s method, randomization tests with 10000 iterations are used to estimate the overall $p$-values. The values of $P_{L}$, $P_{R}$, $Q$ and $p$-values derived at the three frames are shown in Table 6. $H_{0}$ is rejected at all frames when we use Fisher’s method. For Pearson’s method, $H_{0}$ is rejected at the first and last frames. Comparing with Table 4 and 5, the $p$-values for the Fisher’s method are more extreme. This indicates that Fisher’s method is more sensitive to the departures of cleft group from control group.
Table 6: The results of $P_{R}$ and $p$-values obtained from Fisher’s method and Pearson’s method at all three frames using the Smile Data 1.
$P_{L}$ $P_{R}$ $Q$ Fisher’s method Pearson’s method
First frame 206.65 12.26 206.65 0(***) $4\times 10^{-4}$(***)
Middle frame 127.19 33.90 127.19 $3.51\times 10^{-5}$(***) 0.067
Last frame 134.21 26.89 134.21 $6.18\times 10^{-6}$(***) 0.027(*)
5.2.3 Welch’s t-test and Hotelling’s $T^{2}$ Test
There are a few other tests we have applied and could be relevant for some future applications
  • (a)
    The Welch version of the $t$-test (Welch (1947), Lehmann and Romano (2005)), which accommodates different variances for the two groups. In this example, the results of the Welch $t$-test is found to be similar to the standard $t$-test (Table 4). A possible reason is that our sample size is small, so it results in similar outcomes.
  • (b)
@@ 5.2.3 Welch’s t-test and Hotelling’s $T^{2}$ Test
    Hotelling’s $T^{2}$ test (Mardia et al., 2024). We apply here Hotelling’s $T^{2}$ test on unsigned elementary feature vector $\bm{a}_{n}$, $n=1,\ldots,N$. In this example, we reject $H_{0}$ for the first two frames at $5\%$ significance level. The reason could be that a Hotelling’s $T^{2}$ test treats the features separately so it does not gain any power if the features point in the same direction; one-sided Hotelling’s $T^{2}$ could be an appropriate approach, which we will pursue in future. In this example of Smile Data 1, the sample size is less than the dimension of $\bm{a}_{n}$, i.e. $N<J$, so the covariance matrix is singular, which requires some modifications and will be dealt in future.
5.3 Visual Study Related to Three Composite Statistics
(a) First frame
(b) Middle frame
(c) Last frame
Figure 5: Plots for three composite asymmetry scores. From top to bottom, we have $\phi^{*}_{L_{1}}(\bm{a})$ (3.9) (in solid dots), $\phi_{L_{1}}(\bm{a})$ (3.6) (in triangles) and $\phi_{L_{2}}(\bm{a})$ (3.7) (in crosses) at the three frames computed on the Smile Data 1. Plot (a) is at first frame, while (b) is at middle frame, whereas (c) is at last frame. In plot (a), $\phi_{L_{1}}(\bm{a})$ and $\phi_{L_{2}}(\bm{a})$ are divided by 2. In plot (b) and (c), $\phi_{L_{2}}(\bm{a})$ is divided by 3. The first, third and fifth rows from top to bottom are cleft while others are control.
Since the test results indicate significant differences between the two groups, we would like to visualize the distributions of $\phi^{*}_{L_{1}}(\bm{a})$, $\phi_{L_{1}}(\bm{a})$ and $\phi_{L_{2}}(\bm{a})$ defined in (3.9), (3.6) and (3.7) respectively between the two groups. Hence, dot plots similar to the stem leaf plot are created and given in Figure 5. In these dot plots, composite asymmetry scores for both groups are plotted together at each frame. According to these dot plots, substantial overlaps between cleft and control subjects can be seen. Moreover, the three scores used in this section all have similar performances, as the separation between the two groups at each frame revealed by the dot plots are all similar.
Figure 6 shows the dot plots for weighted $\phi_{L_{1}}(\bm{a})$ and weighted $\phi_{L_{2}}(\bm{a})$ simultaneously. According to these figures, these two composite scores push outliers farther away from the rest of the samples, more so for weighted $\phi_{L_{1}}$.
(a) First frame
(b) Middle frame
(c) Last frame
Figure 6: Plots of two weighted asymmetry scores. From top to bottom, we have weighted $\phi_{L_{1}}(\bm{a})$ (in solid dots) and $\phi_{L_{2}}(\bm{a})$ (in triangles) at the three frames computed using the Smile Data 1. The first and third rows from top to bottom are cleft while others are control. Plot (a) is at first frame, while (b) is at middle frame, whereas (c) is at last frame.
5.4 Summary of the Analyses
@@
We end this section by summarizing the main conclusions for these cleft lip data:
  • (a)
@@
    There are statistically significant differences between the two groups at all frames. The control subjects are overall more symmetric than the cleft lip subjects.
  • (b)
@@
    Substantial overlaps can be found on dot plots in Figure 5. This matches the expert’s expectation that some cleft lip subjects should be fairly close to normal subjects after surgeries.
  • (c)